if y = t^3 , x = t^2 . Then find the vale of d^2y / dx^2 = ?

  if   y =  t^3 ,   x  =  t^2 . Then find the vale of d^2y / dx^2 = ?

Grade:11

2 Answers

Khimraj
5 years ago
y = tthen
Khimraj
5 years ago
y =t3
d2y/dt2 = 6t
x = t2
d2x/dt2 = 2
d2y/dx2 = (dy2/dt2)/(dx2/dt2) = 6t/2 = 3t.
Hope it clears.